# Diffence: Fencing Membership Privacy With Diffusion Models

### v1

@@ Abstract
Deep learning models, while achieving remarkable performance across various tasks, are vulnerable to member inference attacks, wherein adversaries identify if a specific data point was part of a model’s training set. This susceptibility raises substantial privacy concerns, especially when models are trained on sensitive datasets. Current defense methods often struggle to provide robust protection without hurting model utility, and they often require retraining the model or using extra data. In this work, we introduce a novel defense framework against membership attacks by leveraging generative models. The key intuition of our defense is to remove the differences between member and non-member inputs which can be used to perform membership attacks, by re-generating input samples before feeding them to the target model. Therefore, our defense works pre-inference, which is unlike prior defenses that are either training-time (modify the model) or post-inference time (modify the model’s output).
A unique feature of our defense is that it works on input samples only, without modifying the training or inference phase of the target model. Therefore, it can be cascaded with other defense mechanisms as we demonstrate through experiments. Through extensive experimentation, we show that our approach can serve as a robust plug-n-play defense mechanism, enhancing membership privacy without compromising model utility in both baseline and defended settings. For example, our method enhanced the effectiveness of recent state-of-the-art defenses, reducing attack accuracy by an average of 5.7% to 12.4% across three datasets, without any impact on the model’s accuracy. By integrating our method with prior defenses, we achieve new state-of-the-art performance in the privacy-utility trade-off.
1 Introduction
Deep learning has achieved remarkable achievements in recent years. However, it’s been demonstrated that deep learning models can memorize information from their training [1, 2, 3], making them susceptible to membership inference attacks (MIAs). Such attacks seek to infer if a specific data point was part of a model’s training set [4]. Given that deep learning models are often trained on sensitive data, including facial images and medical records, the potential success of MIAs poses a significant threat to individual privacy [4].
Figure 1: Examples of original samples and their reconstructions on CIFAR10. Original samples are successfully identified as members by the adversary, while the reconstructed samples are classified as non-members.
Figure 2: An illustration of different defense stages and our proposed defense. Defenses are categorized according to their implementation stages in the machine learning pipeline: the training phase, pre-inference phase, and post-inference phase. Different defense strategies can be deployed at various stages for an integrated defense approach. Our method, uniquely positioned in the pre-inference phase, is compatible with all other methods.
@@ 1 Introduction
Technique
Requires Re-training
Requires Additional Data
Impact on Model Accuracy
Deployment Stage
AdvReg [5]
High
Training
MemGuard [6]
None
Post-Inference
DPSGD [7]
High
Training
SELENA [8]
Low
Training
RelaxLoss [9]
None
Training
HAMP [10]
Low
Training
Ours (Scenario 1)
None
Pre-inference
Ours (Scenario 2)
None
Pre-inference
Ours (Scenario 3)
None
Pre-inference
@@
The growing threat of MIAs has motivated the development of different defense strategies. Existing defense mechanisms can be categorized into two main categories based on their application phase in the machine learning pipeline. The first category encompasses defenses that operate during the training phase. These methods employ privacy-preserving techniques to train models in a way that reduces their memorization of training data, thereby mitigating privacy risks [7, 5, 8, 9, 10]. Key approaches within this category include Differential Privacy Stochastic Gradient Descent (DPSGD) [7], adversarial regularization [5], and knowledge distillation [8]. The second category targets the post-inference phase, focusing on reducing membership privacy leakage by directly rectifying the disparities in the model’s outputs between members and non-members. A representative defense in this category is MemGuard [6], which counters the MIAs by altering the confidence vector to fool the adversary’s attack model. Despite the different defense approaches proposed, existing defenses still suffer from several limitations, as exemplified in Table I. For example, they often struggle to achieve an optimal privacy-utility trade-off [9, 8]. Additionally, some defenses often rely on extra data to support their methods [5, 6].
@@
A new category for MIA defense: In this paper, we introduce a third category of defense mechanisms that operates in the pre-inference stage. A pre-inference defense does not alter model outputs or the model itself, as with the two main categories introduced above. Instead, a pre-inference defense mechanism modifies input samples, before they are sent to the target model for inference. To our best knowledge, no prior work has developed a pre-inference defense against MIAs. Figure 2 compares the three categories of MIA defense techniques, and Table I shows example mechanisms across these categories.
@@
Our contributions. In this paper, we introduce a novel diffusion model-based defense that can enhance the privacy of both the model without defense and those equipped with other defenses, without sacrificing any utility. Advanced generative models such as GANs [11] and diffusion models [12] have been previously employed during inference to safeguard deep learning models from adversarial attacks [13]. However, the application of generative models during inference to defend privacy attacks, specifically MIA, remains largely unexplored. In this paper, we show that diffusion models can also serve as powerful tools integrated into the defense framework against MIAs, preserving membership privacy without harming model utility.
@@
The target model’s memorization of members results in distinct behaviors when classifying seen (member) versus unseen (non-member) samples. MIAs exploit these behavioral discrepancies to distinguish between members and non-members, though different attack methods may leverage different features. In this work, we focus on black-box attacks, which typically infer membership by utilizing information from the sample outputs. He et al. [14] categorized the information exploited in these attacks into two parts: prediction posteriors and labels. We contend that the fundamental cause of MIAs is the non-negligible gap between the model’s outputs for members and non-members. Building on this, we divide the features exploitable by attacks into two categories: the train-to-test accuracy gap and the prediction distribution gap. These features, as utilized in existing attacks [4, 15, 16, 5, 17], are summarized in Table II. Note that some attacks which rely solely on labels [18, 19] may also indirectly exploit the prediction distribution gap through the robustness of the predicted labels.
@@
TABLE II: Summary of the features and information utilized by various attacks. ✓means the specific gap is exploited by the attack, - otherwise. Regarding features, P denotes the use of prediction posteriors and L denotes the use of label.
Attacks
Train-to-test Accuracy Gap
Prediction Distribution Gap
Features
NN-based attacks [15, 4, 5]
P,L
Metric-corr [16]
L
Metric-loss [16]
P,L
Metric-conf [15]
P
Metric-ent [17]
P
Metric-ment [15]
P,L
Label-only attacks [18, 19]
P,L
@@
A successful MIA defense should fundamentally eliminate/shrink the two gaps described above between members and nonmembers. Our defense mainly focuses on eliminating the prediction distribution gap while leaving the samples’ prediction labels unchanged to avoid any loss in utility. This is achieved by reconstructing the samples using a generative model before they are input into the target model. Through sample reconstruction, the model encounters samples in the inference stage that are not exact replicas of those observed during training, regardless of whether they are member or non-member samples. We show that this method effectively reduces the discrepancy in the prediction distributions between members and non-members.
Our method involves generating multiple reconstructed images and selecting the best one that suits our purpose. Among the generated samples, only those with the same predicted label as the original sample are considered as candidates for selection. This ensures that the model’s predictive outcomes remain unchanged, thereby maintaining the model’s accuracy. We designed different sample selection strategies for three scenarios based on the assumptions in the defender. These scenarios include a defender having access to both some member and non-member samples, having access only to members, and having access solely to a trained model without any knowledge of samples’ memberships. We demonstrate that our method is effective across all three scenarios. We show that with more information, defenders can perform a more nuanced analysis of the prediction distribution. This enables a more strategic selection of reconstructed images, forcing the prediction distributions of members and non-members to align more closely, thereby providing robust protection. Figure 1 provides some example inputs and their selected reconstructions. The membership of the original samples was successfully inferred by attackers while their reconstructed versions prevent the disclosure of membership.
While our defense can be implemented using arbitrary generative models, we use diffusion models [20] due to their state-of-the-art performances in generative tasks [12]; we therefore call our mechanism Diffence, i.e., a diffusion-based fence against MIAs. A diffusion model operates in two phases: (i) the forward diffusion phase incrementally adds noise to samples, and (ii) the reverse generation phase generates new samples by applying multi-step denoising. To generate a new sample from an existing one, the original sample is subjected to a specific number of forward diffusion steps, followed by its reverse denoising steps. The procedure is suitable for reconstructing an existing sample in detail. The resultant sample, although distinct from the original in irrelevant details, retains identical semantic attributes, which aligns with our desired defensive objective. Note that our pre-inference defence can be cascaded with other (i.e., training or post-inference) defense mechanisms, due to its minimal deployment constraints and its plug-and-play flexibility.
We conduct comprehensive experiments, comparing our defense mechanisms against six state-of-the-art defense strategies across three datasets using two popular model architectures, ResNet [21] and DenseNet [22]. Our empirical results validate our approach’s effectiveness in enhancing membership privacy without decreasing the model utility, irrespective of the model’s operational context—be it a baseline (vanilla) or a defended setting.
We summarize our contributions as follows:
@@
    We propose a novel diffusion model-based membership inference defense framework, which can enhance the membership privacy of pre-existing (both undefended and defended) models without compromising the utility of the model.
@@
    We propose a new defense pipeline, which for the first time combines defenses deployed at different stages to achieves better defense performance.
@@
    We implemented the prototype of the proposed method. Our extensive experiments show that our proposed defense can effectively improve membership privacy of existing models without utility loss. For example, We successfully reduced the attack accuracy of the most recent state-of-the-art defenses, HAMP and RelaxLoss, from 57.41% to 53.61% and from 63.51% to 60.13% on CIFAR100, which achieved new state-of-the-art performance, without any loss in model accuracy. Furthermore, we show that in certain settings our method can enhance both the utility and privacy of the model.
2 Background and Preliminaries
2.1 Diffusion Models
@@
2.2 Threat Model
2.2.1 Attacker
Black-box Access: Following prior works [6, 10], we assume the attacker has black-box access to the target model. This means the attacker can query the target model using a black-box API and receive corresponding prediction vectors, while direct access to the model’s parameters remains restricted.
Partial Knowledge of Membership: Like previous defenses [8, 10], we assume a strong attacker that knows a limited number of samples from both the training and test sets, i.e., it knows the membership of some members and some non-members. The attack’s goal is to infer the membership of any other unexposed sample.
Full Knowledge of Defense Technique: We assume the attacker is fully aware of the defense technique’s deployment and the architecture of the target model. As a result, for attacks that need shadow models, the attacker can train shadow models that mirror the target model’s training method. In the context of our approach, we assume the attacker has full knowledge of the diffusion model and how it is deployed.
2.2.2 Defender
@@
3 Methodology
Our main idea is to eliminate the differences in predictive behavior between members and non-members, thereby fundamentally mitigating membership privacy leakage. As described in Section 1, we categorize these disparities into the train-to-test accuracy gap and the prediction distribution gap. Our method primarily targets the elimination of the prediction distribution gap. However, by combining our approach with other defenses, we can simultaneously address both gaps, achieving state-of-the-art defense performance. In this section, we first explain how differences in predictive behavior can lead to privacy leakage. We then discuss how we design our defense to narrow these gaps.
3.1 Prediction Gaps between Members and Non-members
Prior studies on MIAs have highlighted a distinct model behavior concerning training and test data, pinpointing it as a key contributor to membership privacy leakage [29]. The gap in prediction distributions between members and non-members can be reflected in various features. These may include assigning higher confidence levels, lower loss, and reduced prediction entropy to members. The disparities in these features are key factors leading to membership privacy leakage, with almost all effective MIAs exploiting one or several of these feature discrepancies [16, 4, 17, 29]. Moreover, these features are interrelated; for instance, lower loss often implies reduced prediction entropy and higher confidence levels. Carlini et al. [29] also proposed parametric modeling of prediction confidence to achieve a more Gaussian-like distribution. Such a distribution more effectively distinguishes the prediction distribution gap between members and non-members. In our subsequent discussions, we adopt this parametric modeling approach to represent the prediction gap. The only difference in our approach is the selection of the maximum confidence from the output vector, rather than the confidence of the correct class, as our focus is on the prediction distribution gap rather than the train-to-test accuracy gap. The parametric modeling function is shown in equation 2.
$\phi(p)=\log(\frac{p}{1-p}),\text{ for }p=max(f(x))$ (2)

### v2

@@ Abstract
Deep learning models, while achieving remarkable performances across various tasks, are vulnerable to membership inference attacks (MIAs), wherein adversaries identify if a specific data point was part of the model’s training set. This susceptibility raises substantial privacy concerns, especially when models are trained on sensitive datasets. Although various defenses have been proposed, there is still substantial room for improvement in the privacy-utility trade-off. In this work, we introduce a novel defense framework against MIAs by leveraging generative models. The key intuition of our defense is to remove the differences between member and non-member inputs, which is exploited by MIAs, by re-generating input samples before feeding them to the target model. Therefore, our defense, called Diffence, works pre inference, which is unlike prior defenses that are either training-time (modify the model) or post-inference time (modify the model’s output).
A unique feature of Diffence is that it works on input samples only, without modifying the training or inference phase of the target model. Therefore, it can be cascaded with other defense mechanisms as we demonstrate through experiments. Diffence is specifically designed to preserve the model’s prediction labels for each sample, thereby not affecting accuracy. Furthermore, we have empirically demonstrated that it does not reduce the usefulness of the confidence vectors. Through extensive experimentation, we show that Diffence can serve as a robust plug-n-play defense mechanism, enhancing membership privacy without compromising model utility—both in terms of accuracy and the usefulness of confidence vectors—across standard and defended settings. For instance, Diffence reduces MIA attack accuracy against an undefended model by 15.8% and attack AUC by 14.0% on average across three datasets, all without impacting model utility. By integrating Diffence with prior defenses, we can achieve new state-of-the-art performances in the privacy-utility trade-off. For example, when combined with the state-of-the-art SELENA defense it reduces attack accuracy by 9.3%, and attack AUC by 10.0%. Diffence achieves this by imposing a negligible computation overhead, adding only 57ms to the inference time per sample processed on average.
I Introduction
Figure 1: Examples of original samples and their reconstructions on CIFAR-10. Original samples are successfully identified as members by the adversary, while the reconstructed samples are classified as non-members.
Figure 2: An illustration of different MIA defense stages and our proposed Diffence. Defenses are categorized according to their implementation stages in the machine learning pipeline: the training phase, pre-inference phase, and post-inference phase. Different defense strategies can be deployed at various stages for an integrated defense approach. Diffence uniquely positioned in the pre-inference phase, is compatible with all other methods.
@@ I Introduction
Technique Requires Re-training Requires Additional Data Impact on Model Accuracy Deployment Stage
AdvReg [1] High Training
MemGuard [2] - None Post-Inference
DPSGD [3] - High Training
SELENA [4] - Low Training
RelaxLoss [5] - None Training
HAMP [6] - Low Training
Diffence (Scenario 1) - None Pre-inference
Diffence (Scenario 2) - - None Pre-inference
Diffence (Scenario 3) - - None Pre-inference
Deep learning has made remarkable achievements in recent years. However, it has been demonstrated that deep learning models can memorize information from their training [7, 8, 9], making them susceptible to membership inference attacks (MIAs). Such attacks seek to infer if a specific data point was part of a model’s training set [10]. Given that deep learning models are often trained on sensitive data, such as facial images [11] and medical records [12], the potential success of MIAs poses a significant threat to privacy [10].
@@
The growing threat of MIAs has motivated the development of different defense strategies. Existing defense mechanisms can be categorized into two main classes based on their application phase in the machine learning pipeline. The first category includes defenses that operate during the training phase. These methods employ privacy-preserving techniques to train models in a way that reduces their memorization of training data, thereby mitigating privacy risks [3, 1, 4, 5, 6]. Key approaches within this category include Differential Privacy Stochastic Gradient Descent (DPSGD) [3], adversarial regularization (AdvReg) [1], and knowledge distillation [4]. The second category targets the post-inference phase, focusing on reducing membership privacy leakage by directly rectifying the disparities in the model’s outputs between members and non-members. A representative defense in this category is MemGuard [2], which modifies output confidence vectors to fool the adversary’s attack model.
@@
A new class of MIA defenses: In this paper, we introduce a third category of defense mechanisms that operate in the pre-inference stage. A pre-inference defense does not alter model outputs or the model itself, as with the two main categories introduced above. Instead, a pre-inference defense mechanism modifies input samples before they are sent to the target model for inference. To our best knowledge, no prior work has developed a pre-inference defense against MIAs. Figure 2 compares the three categories of MIA defense techniques, and Table I shows example mechanisms across these categories.
@@
We introduce Diffence, a novel pre-inference MIA defense that enhances privacy without sacrificing model accuracy. Diffence uses diffusion models. Advanced generative models such as GANs [13] and diffusion models [14] have been previously employed during inference to safeguard deep learning models from adversarial attacks [15]. However, the application of generative models during inference to defend against privacy attacks, specifically MIA, remains largely unexplored. In this paper, we show that diffusion models can also serve as powerful tools integrated into the defense framework against MIAs, preserving membership privacy without harming model utility.
@@
Intuitions behind Diffence: The target model’s memorization of members results in distinct behaviors when classifying seen (member) versus unseen (non-member) samples. MIAs exploit these behavioral discrepancies to distinguish between members and non-members, though different attack methods may leverage different features. In this work, we focus on black-box attacks, which typically infer membership by utilizing information from the sample outputs. He et al. [16] categorized the information exploited in these attacks into two parts: prediction posteriors and labels. We contend that the fundamental cause of MIAs is the non-negligible gap between the model’s outputs for members and non-members. Building on this, we divide the features exploitable by attacks into two categories: the train-to-test accuracy gap and the prediction distribution gap. These features, as utilized in existing attacks [10, 17, 18, 1, 19], are summarized in Table II. Note that some attacks which rely solely on labels [20, 21] may also indirectly exploit the prediction distribution gap through the robustness of the predicted labels.
@@
TABLE II: Summary of the features and information utilized by various MIAs. ✓means the specific gap is exploited by the attack, - otherwise. Regarding features, P denotes the use of prediction posteriors and L denotes the use of label.
Attacks Train-to-test Accuracy Gap Prediction Distribution Gap Features
NN-based attacks [17, 10, 1] P,L
Metric-corr [18] - L
Metric-loss [18] P,L
Metric-conf [17] - P
Metric-ent [19] - P
Metric-ment [17] P,L
Label-only attacks [20, 21] P,L
@@
A successful MIA defense should fundamentally eliminate/shrink the two gaps described above between members and nonmembers while preserving the model’s utility. Diffence primarily focuses on minimizing the prediction distribution gap without altering the samples’ prediction labels, thus maintaining model accuracy. This is achieved by reconstructing the samples using a generative model before they are input into the target model. Through sample reconstruction, the model encounters samples in the inference stage that are not exact replicas of those observed during training, regardless of whether they are member or non-member samples, effectively reducing the discrepancy in the prediction distributions. Meanwhile, our reconstruction process retains the semantic information of the original samples, differing only in details, which allows the model’s output confidence scores to accurately reflect the characteristics of the original sample. Diffence includes generating multiple reconstructions for each sample and implementing various strategic selection techniques based on different assumptions about the defender. These techniques aim to align the prediction distributions of members and non-members as closely as possible while maintaining the prediction label, further details are provided in Section 4.
Diffence’s implementation: While our defense can be implemented using arbitrary generative models, we use diffusion models [22] due to their state-of-the-art performances in generative tasks [14] and compatibility with our defense strategy; we therefore call our mechanism Diffence, i.e., a diffusion-based fence against MIAs. A diffusion model can be used to modify an original sample through forward diffusion and reverse denoising steps, creating a new sample that maintains the same semantic attributes while altering non-essential details, aligning with our defense objectives (details are discussed in Section III-B). Figure 1 provides some example inputs and their selected reconstructions. The diffusion model employed for defense does not need to be exclusively designed for this purpose; hence, defenders can utilize off-the-shelf diffusion models. Additionally, our results demonstrate that defenders can employ diffusion models pre-trained on different datasets to effectively defend and achieve comparable performance. For instance, a diffusion model trained on ImageNet can successfully provide defense for CIFAR10.
We evaluated Diffence on three benchmark datasets (CIFAR-10 [23], CIFAR-100 [23], and SVHN [24]) and two high-resolution image datasets (CelebA [25] and UTKFace [26]) using four model architectures: ResNet18 [27], DenseNet121 [28], VGG16 [29], and Vision Transformer (ViT) [30]. Our experiments comprehensively cover six state-of-the-art MIAs and six state-of-the-art defenses. Note that Diffence can be viewed as a membership privacy booster that can be cascaded with other (i.e., training or post-inference) defense mechanisms, due to its minimal deployment constraints and its plug-and-play flexibility. This also allows us to combine the strengths of other methods. For instance, Diffence can be combined with defenses that address label-only attacks, enhancing their effectiveness against other stronger attacks. While Diffence does not directly counter label-only threats, this combination creates a comprehensive defense strategy effective against all types of attacks. Our empirical results validate Diffence’s effectiveness in enhancing membership privacy without decreasing the model utility, irrespective of the model’s operational context—be it a baseline (vanilla) or a defended setting. For example, we decreased attack accuracy for SELENA, Relaxloss and HAMP by 9.3%, 12.2%, and 14.4%, and attack AUC by 10.0%, 11.4%, and 14.4%, respectively, on average across three benchmark datasets. Diffence also reduced the true positive rate (TPR) under 0.1% false positive rate (FPR) by 52.8% on average across six tested defenses on the SVHN dataset. All of this was achieved without any loss in model utility, with an average additional inference overhead of only 57ms.
In this context of utility, unlike previous studies on MIAs that focused solely on model accuracy, we recognize model utility encompasses not just accuracy, but also the meaningfulness of the output confidence scores or how well the confidences are calirbated [31, 32]. Through our evaluation, we have also verified that Diffence effectively maintains the meaningfulness of confidence scores, which indicates that it indeed preserves the overall utility.
Summary of contributions: Our contributions are as follows:
@@
    We propose a novel diffusion model-based membership inference defense framework called Diffence, which can boost the membership privacy of pre-existing (both undefended and defended) models without compromising the model utility.
@@
    We propose a new defense pipeline, which for the first time combines defenses deployed at different stages to achieve better defense performance.
@@
    We implemented the prototype of the Diffence. Our extensive experiments show that Diffence can effectively improve the membership privacy of existing models without utility loss. Furthermore, we show that in certain settings Diffence can enhance both the model accuracy and privacy of the model.
II Background and Preliminaries
II-A Diffusion Models
@@
II-B Threat Model
II-B1 Attacker
Black-box Access: Following prior works [2, 6], we assume the attacker has black-box access to the target model. This means the attacker can query the target model using a black-box API and receive corresponding prediction vectors, while direct access to any intermediate results, including the reconstructions from the diffusion model and intermediate outputs from the target model remains restricted. Therefore, we assume there is no privacy leakage through the diffusion model or the intermediate results of the target model.
Partial Knowledge of Membership for Members: Like previous defenses [4, 6], we assume a strong attacker that knows a limited number of samples from both the training and test sets, i.e., they know the membership status of some members and some non-members. Therefore, the attacker can directly train their attack classifier using the known members and non-members of the target model without the need to train any shadow models. The goal of the attack is to infer the membership status of any other unexposed sample.
Full Knowledge of Defense Technique: We assume the attacker is fully aware of the deployment of the defense technique and the architecture of the target model. In the context of Diffence, we assume the attacker has full knowledge of how the diffusion model is deployed. All attacks against Diffence in our experiments can be considered adaptive attacks, as we assume the attacker can process samples in the same manner as Diffence when training their attack classifier. This means the attacker can use reconstructions from the same diffusion model to train the attack classifier instead of using the original samples.
II-B2 Defender
@@
III Methodology
Our main idea is to eliminate the differences in predictive behavior between members and non-members, thereby fundamentally mitigating membership privacy leakage. As described in Section I, we categorize these disparities into the train-to-test accuracy gap and the prediction distribution gap. Diffence primarily targets the elimination of the prediction distribution gap. However, Diffence can seamlessly integrate with other defenses that reduce the train-to-test accuracy gap, such as SELENA [4], enabling a comprehensive defense strategy that simultaneously addresses both gaps without sacrificing the model’s utility, achieving state-of-the-art defense performance. In this section, we first explain how differences in predictive behavior can lead to privacy leakage. We then introduce how we design Diffence to narrow these gaps.
III-A Intuition Behind Diffence
Prior studies on MIAs have highlighted a distinct model behavior concerning training and test data, pinpointing it as a key contributor to membership privacy leakage [39]. The gap in prediction distributions between members and non-members can be reflected in various features. These may include assigning higher confidence levels, lower loss, and reduced prediction entropy to members. The disparities in these features are key factors leading to membership privacy leakage, with almost all effective MIAs exploiting one or several of these feature discrepancies [18, 10, 19, 39]. Moreover, these features are interrelated; for instance, lower loss often implies reduced prediction entropy and higher confidence levels.
Carlini et al. [39] proposed parametric modeling of prediction confidence to achieve a more Gaussian-like distribution. Such a distribution more effectively distinguishes the prediction distribution gap between members and non-members. In our subsequent discussions, we adopt this parametric modeling approach to represent the prediction gap. The only difference is that we use the maximum confidence from the output vector rather than the confidence of the correct class, as our focus is on the prediction distribution gap rather than the train-to-test accuracy gap. The parametric modeling function is shown in Equation 2.
$\phi(p)=\log\left(\frac{p}{1-p}\right),\text{ for }p=\max(f(x))$ (2)
where $f(x)$ is the model’s output vector for input $x$.